the metro transit system (mcts) is made up of the red line and the blue line. the red line has 4 cars and averages 478 passengers per run. the blue line has 20 cars and averages 468 passengers per run.

determine the relative u fairness if the apportionment if the additional car is given to the red line
Answer #1

determine the average number of passengers per car

the average number of passengers per car is expressed as:
=

the average passengers per car of the two lines are as follows:

red line: 478/4 = 119.5

blue line: 468/20 = 23.4

If an additional car is given to red line, its average passengers become:

red line (new) = 478/5 = 95.6

therefore, red line will have the smaller average passengers per car.

The relative unfairness of apportionment: